# HV-BEV: Decoupling Horizontal and Vertical Feature Sampling for Multi-View 3D Object Detection

### v2

@@ Abstract
The application of vision-based multi-view environmental perception system has been increasingly recognized in autonomous driving technology, especially the BEV-based models. Current state-of-the-art solutions primarily encode image features from each camera view into the BEV space through explicit or implicit depth prediction. However, these methods often focus on improving the accuracy of projecting 2D features into corresponding depth regions, while overlooking the highly structured information of real-world objects and the varying height distributions of objects across different scenes. In this work, we propose HV-BEV, a novel approach that decouples feature sampling in the BEV grid queries paradigm into horizontal feature aggregation and vertical adaptive height-aware reference point sampling, aiming to improve both the aggregation of objects’ complete information and generalization to diverse road environments. Specifically, we construct a learnable graph structure in the horizontal plane aligned with the ground for 3D reference points, reinforcing the association of the same instance across different BEV grids, especially when the instance spans multiple image views around the vehicle. Additionally, instead of relying on uniform sampling within a fixed height range, we introduce a height-aware module that incorporates historical information, enabling the reference points to adaptively focus on the varying heights at which objects appear in different scenes. Extensive experiments validate the effectiveness of our proposed method, demonstrating its superior performance over the baseline across the nuScenes dataset. Moreover, our best-performing model achieves a remarkable 50.5% mAP and 59.8% NDS on the nuScenes testing set.
@@ I Introduction
Figure 1: The multi-view feature sampling comparison between (a) BEVFormer and (b) our HV-BEV. BEVFormer extracts local area features by uniformly sampling several reference points across the global height range within each pillar-like query and projecting them onto corresponding views. In contrast, our HV-BEV adopts adaptive height-aware sampling and constructs a graph for each reference point on the horizontal BEV plane, aggregating features from relevant regions to enhance feature coherence and improve network adaptability.
@@
3D object detection is a fundamental task in intelligent vehicle perception systems, focused on estimating the locations, sizes, and categories of key objects in traffic scenarios—such as vehicles, pedestrians, cyclists, and traffic cones—in the ego-vehicle coordinate system [1]. Although LiDAR provides direct distance measurements that facilitate 3D detection [2, 3, 4, 5, 6], its high cost and limited maintainability make cameras—cost-effective devices capable of capturing rich semantic information—an essential choice for deployment. By strategically positioning cameras at various angles around the vehicle, the system can achieve comprehensive perception of its surroundings. Consequently, multi-view visual 3D object detection remains one of the central focuses of ongoing research [7, 8, 9].
@@
Recently, BEV-based 3D perception has gained widespread attention [11, 12, 10] due to the ability of BEV feature representation to model a complete, scale-consistent environment around the vehicle within a unified coordinate system, making it more suitable for downstream tasks such as path planning and control. BEV-based 3D object detection can be broadly categorized into two major methods: explicit and implicit depth estimation [8], based on how 2D features are transformed into 3D space. Explicit depth-based methods primarily utilize depth estimation networks to predict depth distribution in perspective views (PV), transforming PV feature maps into 3D space using determined camera parameters [13, 14, 15, 16]. Instead of the intermediate depth estimation step, implicit methods rely on sophisticated view transformation networks to extract 3D information embedded in the 2D images without the need for an depth map [17, 18, 19, 21, 22].
@@
Most of these existing methods focus on minimizing the feature localization errors caused by depth estimation inaccuracies, while neglecting the 3D structural information of real-world objects. In other words, they pay little attention to the global representation of regions associated with an object when encoding BEV features, resulting in insufficient feature aggregation. While some works have addressed this issue with promising results [23, 24, 25], they are predominantly centered around sparse object query-based schemes, which are specifically designed for the single task of 3D object detection and are difficult to extend to other tasks, i.e., map segmentation, by simply replacing the task head or transform into multi-task models. In addition, most BEV-based methods either ”splat” frustum features through Voxel/Pillar Pooling [27, 13, 15] or directly cross-encode image features into a flattened BEV grid plane [18], leading to the collapse of the height dimension and causing the model to overlook crucial height information. Different types of objects may exist at various heights [26], and we believe this height information should be fully utilized to assist the network in exploring the intrinsic properties of different objects. Moreover, the network should adaptively focus on the potential local heights of objects in different scenes, without dispersing its attention across all possible heights within the global height range. This is because, as we know, in real-world road scenarios, it is rare to encounter multiple objects occupying different heights at the same location simultaneously.
@@
In this paper, we take BEVFormer [18], a representative work of the BEV grid query-based approach, as our baseline and propose a comprehensive 3D object detection framework named HV-BEV that decouples multi-view feature sampling in the horizontal BEV plane and vertical height direction. In the vanilla BEVFormer, multi-view feature sampling is achieved by projecting the 3D reference points located at the center of the corresponding cell region of each BEV query and uniformly sampled across the global height range to the hit views and applying deformable attention [20], as shown in Fig. 1(a). This sampling approach retrieves the corresponding image features for each BEV grid region from different views and implicitly modeling height information; However, it does not fully account for the structured features of 3D objects, particularly the cohesive nature of height information and the interdependencies across different dimensions. We identified that large or near-ego objects are sometimes missed or inaccurately estimated due to occupying multiple BEV grid regions or appearing in multiple camera views through observation. This phenomenon is primarily caused by insufficient feature aggregation in the horizontal dimension. To address this, we propose a horizontal cross-view feature aggregation module, which constructs a dynamic graph for each 3D reference point on the horizontal BEV plane to aggregate global features of the object from different grid regions and views, seen in Fig. 1(b). However, relying solely on horizontal feature aggregation is not enough, as different objects do not exist along a single horizontal line; rather, their heights vary with scene context and object class. Motivated by this, we introduce an adaptive height-aware sampling strategy as shown in Fig. 1(b). The height distributions from aligned historical BEV features and current BEV queries are fused and guide vertical height sampling of 3D reference points, allowing the network to adaptively focus on the most relevant height for each object. The training of this module is done in a sparse way, using only 3D annotated bounding boxes without extra LiDAR points supervision. By decoupling feature sampling in different dimensions, the model can concentrate on information of value across various directions, yielding a more efficient global feature sampling strategy. Notably, while this paper evaluates the performance of HV-BEV using 3D object detection as the primary task for model training and prediction, our approach can be easily extended to other BEV-based perception tasks, such as map segmentation and lane detection.
@@
  • We propose a multi-view 3D object detection method based on the BEV grid query paradigm, which decouples horizontal and vertical feature sampling to fully exploit 3D spatial information, achieving superior detection performance.
  • We design a horizontal cross-view feature aggregation module that fuses relevant structured features of the object from different camera images and grid cell regions.
  • We propose an adaptive height-aware sampling strategy in the vertical dimension to guide the network in identifying relevant object heights of interest. To achieve this, we introduce a deep supervision mechanism that operates without requiring additional LiDAR information.
  • We conduct extensive experiments on the challenging public dataset nuScenes to validate the effectiveness of the proposed method. Our HV-BEV outperforms the baseline model across all evaluated metrics and exhibits outstanding performance within the camera-based 3D object detection community.
@@ III-A Revisiting to BEVFormer-Style Feature Sampling
where $N_{\text{ref}}$ denotes the number of 3D reference points uniformly sampled within a predetermined height range for each query $Q_{p}$, centered at the horizontal coordinate $p(x,y)$ and $\mathcal{V}_{hit}$ represents the hit views associated with the $j$-th reference point, obtained by projecting it via the function $\mathcal{P}(p,i,j)$. Following the temporal and spatial feature interactions via TSA and SCA, the resulting BEV queries are subjected to passed through an Add&Norm layer, then a forward propagation network (FFN) [34] and the last Add&Norm layer, yielding the final refined BEV features $B_{t}$.
@@ III-B Overall Framework
Figure 2: Overall architecture of HV-BEV. Aligning with BEVFormer, we begin by using an image encoder to extract multi-view image features. Within the six-layer BEV encoder, we introduce two decoupled modules that respectively handle intrinsic height awareness and cross-view feature aggregation. The Vertical Adaptive Height-Aware module is designed to capture the potential height distribution of objects within each BEV grid, aiding in the sampling of 3D reference points. Leveraging these height-aware 3D reference points, the Horizontal Cross-View Feature Aggregation module constructs a graph structure on the BEV plane to explore spatial features from neighboring regions.
As illustrated in Fig. 2, our HV-BEV follows the 6-layer encoder architecture of BEVFormer, with the primary difference being an improved multi-view feature sampling method within each layer. Specifically, we first employ an image encoder—comprising a backbone, such as ResNet [44], and a neck, like FPN [45]—to extract multi-view features $F_{t}$ from the multi-view images at time $t$, which then serve as input for the horizontal cross-view feature aggregation module in the next BEV encoder.
A BEV encoder layer receives learnable BEV queries $Q$ and the preserved BEV features $B_{t-1}$ from the previous time step as input. We first apply a standard temporal self-attention layer to query temporal information from historical BEV as Eq.(1). Following this, a vertical height-aware sampling module (Sec. III-C) integrates the predicted height distribution from each $Q_{p}$ of the current BEV queries, obtained through a shared MLP layer, with the historical height distribution generated from $B_{t-1}$ to sample $N_{ref}$ discrete height values from the fused height distribution, setting these as the heights of the $N_{ref}$ 3D reference points for each query $Q_{p}$. Notably, by introducing deep supervision into this module, the BEV queries are guided to focus on the height ranges where objects are most likely to appear. In the horizontal cross-view feature aggregation module (Sec. III-D), each 3D reference point learns a dynamic graph within its height-specific BEV plane, projecting all the nodes in this graph onto multi-view images to fully aggregate spatial features from hit views. This approach allows for comprehensive integration of neighboring regions’ spatial information that may appear in different views, further refining BEV features. After passing through additional standard layers (FFN, Add&Norm), the refined BEV features are fed as input queries to the successive BEV encoder layer and iteratively enhanced, producing the final BEV features, which are then utilized by task-specific heads for 3D detection and other perception tasks.
@@ III-C Vertical Adaptive Height-Aware Sampling
In HeightFormer, the authors use attention weights to compute a weighted average of predefined anchor heights within a BEV sample, thereby visualizing how common SCA implicitly encodes height information. Instead of sampling at fixed heights across the entire range, we propose an adaptive height-aware sampling method that enables the model to dynamically adjust the optimal sampling heights based on the specific scene characteristics.
Through TSA, each query $Q_{p}$ looks up temporal information from its associated regions in history BEV $B_{t-1}$. To incorporate historical height information, we further generate a historical height distribution $H_{p}^{hist}\in\mathbb{R}^{1\times D}$ from $B_{t-1}$, where $D$ denotes the number of discrete height bins over the global height range, by first establishing a mapping $\mathcal{M}$ from time $t$ to $t-1$ enabling us to locate the previous position $p_{t-1}$ of $p$ prior to ego-motion. This mapping is defined as follows:
$\begin{array}[]{l}\mathcal{M}:p_{t-1}\leftarrow f_{t,t-1}^{T}(f_{t,t-1}^{R}(p,\mathbf{R}_{t,t-1}),\mathbf{T}_{t,t-1})\end{array},$ (3)
where $f_{t,t-1}^{T}$ and $f_{t,t-1}^{R}$ denote the translation and rotation functions, $\mathbf{T}_{t,t-1}$ and $\mathbf{R}_{t,t-1}$ represent the translation vector and the rotation matrix of ego-motion from time $t$ to $t-1$ respectively. The relevant feature vector $B_{t-1}^{p}$ associated with the grid region centering at $p_{t-1}$ can be obtained to generate $H_{p}^{hist}$ through the shared MLP and a softmax function, which are also utilized to estimate current height distribution $H_{p}^{cur}\in\mathbb{R}^{1\times D}$ of $Q_{p}$.
@@
Following the above process, each query in $Q$ produces both a current height distribution and a historical height distribution, which are then organized across the BEV space to form complete distributions for the entire BEV plane, represented as $H^{cur},H^{hist}\in\mathbb{R}^{H\times W\times D}$. Unlike depth, the height of a moving object remains constant, allowing us to further refine the current height distribution prediction by merging $H^{cur}$ and $H^{hist}$. Inspired by [26], we introduce a dual-branch cross-attention (DCA) fusion strategy to effectively integrate these height distributions. As shown in Fig.3, it consists of two deformable attention branches. In each branch, the height distribution $H_{p}^{cur/hist}$ at each grid cell serves as the query, aggregating information related to the grid center position $p$ from $H^{hist/cur}$ in another branch:
@@
The output distributions from both branches are then summed and passed through an MLP layer to generate the fused height distribution $H^{fus}\in\mathbb{R}^{H\times W\times D}$. For each resulting height distribution $H_{p}^{fus}$, we apply a top-K function to sample $N_{ref}$ discrete height values, which are subsequently used as the height coordinates of the 3D reference points in the horizontal cross-view feature aggregation.
III-D Horizontal Cross-View Feature Aggregation

### v3

@@ Abstract
The application of vision-based multi-view environmental perception system has been increasingly recognized in autonomous driving technology, especially the BEV-based models. Current state-of-the-art solutions primarily encode image features from each camera view into the BEV space through explicit or implicit depth prediction. However, these methods often overlook the structured correlations among different parts of objects in 3D space and the fact that different categories of objects often occupy distinct local height ranges. For example, trucks appear at higher elevations, whereas traffic cones are near the ground. In this work, we propose a novel approach that decouples feature sampling in the BEV grid queries paradigm into Horizontal feature aggregation and Vertical adaptive height-aware reference point sampling (HV-BEV), aiming to improve both the aggregation of objects’ complete information and awareness of diverse objects’ height distribution. Specifically, a set of relevant neighboring points is dynamically constructed for each 3D reference point on the ground-aligned horizontal plane, enhancing the association of the same instance across different BEV grids, especially when the instance spans multiple image views around the vehicle. Additionally, instead of relying on uniform sampling within a fixed height range, we introduce a height-aware module that incorporates historical information, enabling the reference points to adaptively focus on the varying heights at which objects appear in different scenes. Extensive experiments validate the effectiveness of our proposed method, demonstrating its superior performance over the baseline across the nuScenes dataset. Moreover, our best-performing model achieves a remarkable 50.5% mAP and 59.8% NDS on the nuScenes testing set. The code is available at https://github.com/Uddd821/HV-BEV.
@@ I Introduction
Figure 1: The multi-view feature sampling comparison between (a) BEVFormer and (b) our HV-BEV. BEVFormer extracts local area features by uniformly sampling several reference points across the global height range within each pillar-like query and projecting them onto corresponding views. In contrast, HV-BEV adopts adaptive height-aware sampling and constructs a graph for each reference point on the horizontal plane, aggregating features from relevant regions to enhance feature coherence.
@@
3D object detection is a fundamental task in intelligent vehicle perception systems, focused on estimating the locations, sizes, and categories of key objects in traffic scenarios—such as vehicles, pedestrians, cyclists, and traffic cones—in the ego-vehicle coordinate system [1]. Although LiDAR provides direct distance measurements that facilitate 3D detection [2, 3, 4], its high cost and limited maintainability make cameras—cost-effective devices capable of capturing rich semantic information—an essential choice for deployment. By strategically positioning cameras at various angles around the vehicle, the system can achieve comprehensive perception of its surroundings. Consequently, multi-view visual 3D object detection remains one of the central focuses of ongoing research [5, 6].
@@
Recently, BEV(bird’s eye view)-based 3D perception has gained widespread attention [8, 9, 7] due to the ability of BEV feature representation to model a complete, scale-consistent environment around the vehicle within a unified coordinate system, making it more suitable for downstream tasks such as path planning and control. BEV-based 3D object detection can be broadly categorized into two major methods: explicit and implicit depth estimation [5], based on how 2D features are transformed into 3D space. Explicit depth-based methods primarily utilize depth estimation networks to predict depth distribution in perspective views (PV), transforming PV feature maps into 3D space using determined camera parameters [10, 11, 12, 13]. Instead of the intermediate depth estimation step, implicit methods rely on sophisticated view transformation networks to extract 3D information embedded in the 2D images without the need for an depth map [14, 15, 16, 17, 18].
@@
Most of these existing methods focus on minimizing the feature localization errors caused by depth estimation inaccuracies, while neglecting the 3D structural information of real-world objects. In other words, they pay little attention to the global representation of regions associated with an object when encoding BEV features, resulting in insufficient feature aggregation. While some works have addressed this issue with promising results [20, 21, 22], they are predominantly centered around sparse object query-based schemes, which are specifically designed for the single task of 3D object detection and are difficult to extend to other tasks, i.e., map segmentation, by simply replacing the task head or transform into multi-task models. In addition, most BEV-based methods either ”splat” frustum features through Voxel/Pillar Pooling [24, 10, 12] or directly cross-encode image features into a flattened BEV grid plane [15], leading to the collapse of the height dimension and causing the model to overlook crucial height information. Different object categories tend to occupy distinct local height ranges [23], yet treating all vertical positions equally during feature sampling can lead to ineffective aggregation. For instance, traffic cones typically reside near the ground, but sampling features uniformly along the entire height axis within their corresponding pillar introduces a large amount of irrelevant information, resulting in wasted computation and degraded detection performance. To address this, the network should adaptively focus on the local height regions where objects are likely to appear and, in conjunction with horizontally aggregated features, construct more compact and discriminative 3D representations.
@@
In this paper, we take BEVFormer [15], a representative work of the BEV grid query-based approach, as our baseline and propose a comprehensive 3D object detection framework named HV-BEV that decouples multi-view feature sampling in the horizontal BEV plane and vertical height direction. In the vanilla BEVFormer, multi-view feature sampling is achieved by projecting the 3D reference points located at the center of the corresponding cell region of each BEV query and uniformly sampled across the global height range to the hit views and applying deformable attention [19], as shown in Fig. 1(a). This sampling approach retrieves the corresponding image features for each BEV grid region from different views and implicitly modeling height information; However, it does not fully account for the structured features of 3D objects, particularly the cohesive nature of height information and the interdependencies across different dimensions. We identified that large or near-ego objects are sometimes missed or inaccurately estimated due to occupying multiple BEV grid regions or appearing in multiple camera views through observation. This phenomenon is primarily caused by insufficient feature aggregation in the horizontal dimension. To address this, we propose a dynamic horizontal cross-view feature aggregation (DHCA) module, which constructs a set of neighboring points for each 3D reference point on the horizontal BEV plane to aggregate global features of the object from different grid regions and views, seen in Fig. 1(b). However, relying solely on horizontal feature aggregation is not enough, as different objects do not exist along a single horizontal line; rather, their heights vary with scene context and object class. Motivated by this, we introduce an adaptive height-aware sampling strategy as shown in Fig. 1(b). The height distributions from aligned historical BEV features and current BEV queries are fused and guide vertical height sampling of 3D reference points, allowing the network to adaptively focus on the most relevant height for each object. The training of this module is done in a sparse way, where discrete height distributions are generated based on the centers of 3D annotated bounding boxes to supervise object height estimation. By decoupling feature sampling in different dimensions, the model can concentrate on information of value across various directions, yielding a more efficient global feature sampling strategy. Notably, while this paper evaluates the performance of HV-BEV using 3D object detection as the primary task for model training and prediction, our approach can be easily extended to other BEV-based perception tasks, such as map segmentation and lane detection.
@@
  • We propose a multi-view 3D object detection method based on the BEV grid query paradigm, which decouples horizontal and vertical feature sampling to fully exploit 3D spatial information, achieving superior performance.
  • We design a dynamic horizontal cross-view feature aggregation module that fuses relevant structured features of the object from different camera images and grid cell regions and an adaptive height-aware sampling strategy in the vertical dimension to guide the network in identifying relevant object heights of interest.
  • Extensive experiments conducted on the nuScenes and Lyft datasets to validate the effectiveness of the proposed method. Our HV-BEV outperforms the baseline model across all evaluated metrics and exhibits outstanding performance within the camera-based 3D object detection community.
@@ III-A Revisiting to BEVFormer-Style Feature Sampling
where $N_{\text{ref}}$ denotes the number of 3D reference points uniformly sampled within a predetermined height range for each query $Q_{p}$, centered at the horizontal coordinate $p(x,y)$. $\mathcal{P}(p,i,j)$ represents the projection of the $j$-th reference point in $Q_{p}$ onto the 2D image plane of the $i$-th camera view using the camera intrinsic matrix, and $\mathcal{V}_{hit}$ denotes the set of views in which the reference point is successfully projected. Following the temporal and spatial feature interactions via TSA and SCA, the resulting BEV queries are subjected to passed through an Add&Norm layer, then a forward propagation network (FFN) [31] and the last Add&Norm layer, yielding the final refined BEV features $B_{t}$.
@@ III-B Overall Framework
Figure 2: Overall architecture of HV-BEV. Aligning with BEVFormer, we begin by using an image encoder to extract multi-view image features. Within the six-layer BEV encoder, we introduce two decoupled modules that respectively handle intrinsic height awareness and cross-view feature aggregation. The Vertical Adaptive Height-Aware module is designed to capture the potential height distribution of objects within each BEV grid, aiding in the sampling of 3D reference points. Leveraging these height-aware 3D reference points, the Dynamic Horizontal Cross-View Feature Aggregation module constructs a set of neighboring points on the BEV plane for each one to explore spatial features from neighboring regions.
As illustrated in Fig. 2, our HV-BEV follows the 6-layer encoder architecture of BEVFormer, with the primary difference being an improved multi-view feature sampling method within each layer. Specifically, we first employ an image encoder—comprising a backbone, such as ResNet [41], and a neck, like FPN [42]—to extract multi-view features $F_{t}$ from the multi-view images at time $t$, which then serve as input for the dynamic horizontal cross-view feature aggregation module in the next BEV encoder.
A BEV encoder layer receives learnable BEV queries $Q$ and the preserved BEV features $B_{t-1}$ from the previous time step as input. We first apply a standard temporal self-attention layer to query temporal information from historical BEV as Eq.(1). Following this, a vertical height-aware sampling (VHA) module (Sec. III-C) integrates the predicted height distribution from each $Q_{p}$ of the current BEV queries, obtained through a shared MLP layer, with the historical height distribution generated from $B_{t-1}$. From this fused distribution, the module samples $N_{ref}$ discrete height values using a top-k operation and assigns them as the height coordinates of the $N_{ref}$ 3D reference points corresponding to each query $Q_{p}$. Notably, by introducing deep supervision into this module, the BEV queries are guided to focus on the height ranges where objects are most likely to appear. In the dynamic horizontal cross-view feature aggregation (DHCA) module (Sec. III-D), each 3D reference point learns a set of neighboring points determined by learnable 2D offsets within its height-specific BEV plane. All points in this graph are then projected onto multi-view images to fully aggregate spatial features from the hit views. This approach allows for comprehensive integration of neighboring regions’ spatial information that may appear in different views, further refining BEV features. After passing through additional standard layers (FFN, Add&Norm), the refined BEV features are fed as input queries to the successive BEV encoder layer and iteratively enhanced, producing the final BEV features, which are then utilized by task-specific heads for 3D detection and other perception tasks.
@@ III-C Vertical Adaptive Height-Aware Sampling
In HeightFormer[37], the authors use attention weights to compute a weighted average of predefined anchor heights within a BEV sample, thereby visualizing how common SCA implicitly encodes height information. Instead of sampling at fixed heights across the entire range, we propose an adaptive height-aware sampling method that enables the model to dynamically adjust the optimal sampling heights based on the specific scene characteristics.
Through TSA, the current BEV queries $Q$ can retrieve temporal information from the historical BEV $B_{t-1}$ to generate the fused queries $Q^{\prime}\in\mathbb{R}^{H\times W\times C}$. To further integrate historical height information for predicting the current height distribution, we first employ a shared MLP layer to independently predict the discrete historical and current height distributions from $B_{t-1}$ and $Q^{\prime}$, resulting in $H^{hist},H^{cur}\in\mathbb{R}^{H\times W\times D}$, where $D$ denotes the number of discrete height bins over the global height range. However, due to ego-motion, it is necessary to identify the height distribution at the previous position $p_{t-1}$ corresponding to the grid region centered at the current location $p$. This is achieved by establishing a mapping $\mathcal{M}$ from time $t$ to time $t-1$, defined as follows:
$\begin{array}[]{l}\mathcal{M}:p_{t-1}\leftarrow f_{t,t-1}(p,\mathbf{R}_{t,t-1},\mathbf{T}_{t,t-1})\end{array},$ (3)
$\begin{array}[]{l}f_{t,t-1}(p,\mathbf{R}_{t,t-1},\mathbf{T}_{t,t-1})=\mathbf{R}_{t,t-1}\cdot p+\mathbf{T}_{t,t-1}\end{array},$ (4)
where $\mathbf{R}_{t,t-1}\in\mathbb{R}^{2\times 2}$ and $\mathbf{T}_{t,t-1}\in\mathbb{R}^{2\times 1}$ denote the rotation matrix around the vertical $z$ axis and the horizontal translation vector from time $t$ to $t{-}1$, respectively. Through this mapping, the historical height distribution at $p_{t-1}$ can be warped into the coordinate system at time $t$, resulting in the warped historical height distribution $H^{hist}_{t-1\rightarrow t}\in\mathbb{R}^{H\times W\times D}$. Subsequently, the historical and current height information is further fused.
@@
Unlike depth, the height of a moving object remains constant, allowing us to refine the current height distribution prediction by merging $H^{cur}$ and $H^{hist}_{t-1\rightarrow t}$. Inspired by [23], we introduce a dual-branch cross-attention (DCA) fusion strategy to effectively integrate these height distributions. As shown in Fig.3, it consists of two deformable attention branches. In each branch, the height distribution $H_{p}^{cur/hist}$ at each grid cell serves as the query, aggregating information related to the grid center position $p$ from $H^{hist/cur}$ in another branch (For simplicity, the subscript of $H^{hist}_{t-1\rightarrow t}$ is omitted here.):
@@
The output distributions from both branches are then summed and passed through an MLP layer to generate the fused height distribution $H^{fus}\in\mathbb{R}^{H\times W\times D}$. For each resulting height distribution $H_{p}^{fus}$, we apply a top-k function to sample $N_{ref}$ discrete height values, which are subsequently used as the height coordinates of the 3D reference points in the dynamic horizontal cross-view feature aggregation.
During training, we apply sparse supervision to $H^{fus}$ through utilizing the height of the annotated 3D bounding box center. Specifically, to facilitate the calculation of the discrepancy between the ground-truth height and the discretized predicted height distribution, we first derive the discretized ground-truth height distribution $H_{p}^{gt}$ corresponding to each BEV grid region. Since an object’s height is inherently a range rather than a fixed value, we employ a Gaussian kernel $K$ to compute the differences between the annotated center height of an object and the predefined height bins. Height bins closer to the center are assigned higher probabilities. For convenience, we use an indicator $I_{p}$ to denote whether the cell grid centered at $p$ falls within the object’s region. The probability corresponding to the $m$-th height bin of $H_{p}^{gt}$ is calculated as:
where $z_{gt}$ denotes the height of ground truth center, $z_{m}$ represents the center value of the $m$-th height bin, and $\sigma$ is the standard deviation controlling the spread of the Gaussian kernel. When $I_{p}=0$, assigning equal probability to each height bin encourages the network to attend uniformly to the entire height range, allowing it to handle potentially missed or unannotated objects. The height distribution is supervised as follows:
III-D Dynamic Horizontal Cross-View Feature Aggregation